OpenStudy (anonymous):

how do you simplify i^123

OpenStudy (solomonzelman):

\(\large\color{blue}{ \rm i^{123}=i^{120+3}=i^{120}\times i^3=i^3=? }\)

OpenStudy (anonymous):

method 2 take the integer remainder when you divide \(123\) by \(4\) since \(4\) goes in to \(120\) evenly the remainder is \(3\) so \(i^3\) this is exactly the same as @SolomonZelman method just without writing it down

OpenStudy (cwrw238):

i^3 = i^2 * i = ?

OpenStudy (anonymous):

@cwrw238 person is gone

OpenStudy (cwrw238):

thats the story of my life ! no one listens lol